OpenStudy (anonymous):

5(4+3i)^x+(1-i)^4=5^x solution

OpenStudy (jamesj):

Like the last problem, write everything out in polar form first and then things will become much easier. For example, the first expression on the LHS 5(4 + 3i)^x = 5.(5exp(alpha.i))^x = 5^(x+1).exp(x.alpha.i) where alpha = arctan(3/4)

OpenStudy (anonymous):

can't wait to see this

OpenStudy (jamesj):

Then note that the RHS is real, so the LHS must be as well. So once you've found one expression for the LHS, set the imaginary component to zero.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

no first step is \[5\times 5e^{i\alpha x}-5^x=4\]

OpenStudy (jamesj):

you can see that sin(alpha.x) = 0, and that gives you a big leg up.

OpenStudy (anonymous):

second is \[5^{x+1}e^{i \alpha x}-5^x=4\]?